OpenStudy (anonymous):

Complete the square. 4x^2+24x+29 I completely forgot how to do this.

OpenStudy (anonymous):

(24/2)^2

OpenStudy (ranga):

\[\Large ax ^{2} + bx + c = a(x + d)^{2} + e\] The d can be quickly found by: b/2a

OpenStudy (ranga):

a = 4, b = 24, c = 29 d = b/2a = 24 / 8 = 3 So 4x^2 + 24x + 29 = 4(x + 3)^2 + e find e.

OpenStudy (ranga):

4x^2 + 24x + 29 = 4(x + 3)^2 + e = 4(x^2 + 6x + 9) + e 4x^2 + 24x + 29 = 4x^2 + 24x + 36 + e 29 = 36 + e e = -7 4x^2 + 24x + 29 = 4(x + 3)^2 - 7